OpenStudy (anonymous):

Consider the equation 7x + 3y = 42. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points)

OpenStudy (anonymous):

for part 1....... 7x+3y=42 3y=-7x+42 y=-7/3x +14

OpenStudy (anonymous):

SINCE coefficent of x is negative it makes an obtuse angle with the x-axis and when x=0, y= 14 so the y-intercept is 14..now,